OpenStudy (anonymous):

Consider the equation 4x + 6y = 48. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar.

OpenStudy (anonymous):

4x + 6y = 48. solve for y

OpenStudy (anonymous):

reorder and get 6y=48-4x -> y=-4x/6+8 ->y=-2x/3+8 so slope is -2/3 and y-intersect is 8

OpenStudy (anonymous):

y = -(2/3)*x+8 

OpenStudy (anonymous):

I need the explanation on the graph. Like writing it down in words.

OpenStudy (anonymous):

well whenever you solve something like 4x + 6y = 48. for y you get it in something called y intercept form and it looks like this y=m*x+b Where m is slope and b is the y intercept when x=0 so if you solve you get y = -(2/3)*x+8 it means slope is -2/3 and y intercept is 8

OpenStudy (anonymous):

Now can you explain how I would graph that? In words not a picture. please.